# Project Page: https://yikai-wang.github.io/asuka/ MISATO Dataset: https://github.com/Yikai-Wang/asuka-misato

### v2

@@ Abstract
Recent progress in inpainting increasingly relies on generative models, leveraging their strong generation capabilities for addressing large irregular masks. However, this enhanced generation often introduces context-instability, leading to arbitrary object generation within masked regions. This paper proposes a balanced solution, emphasizing the importance of unmasked regions in guiding inpainting while preserving generation capacity. Our approach, Aligned Stable Inpainting with UnKnown Areas Prior (ASUKA), employs a Masked Auto-Encoder (MAE) to produce reconstruction-based prior. Aligned with the powerful Stable Diffusion inpainting model (SD), ASUKA significantly improves context stability. ASUKA further adopts an inpainting-specialized decoder, highly reducing the color inconsistency issue of SD and thus ensuring more visual-consistent inpainting. We validate effectiveness of inpainting algorithms on benchmark dataset Places 2 and a collection of several existing datasets, dubbed MISATO, across diverse domains and masking scenarios. Results on these benchmark datasets confirm ASUKA’s efficacy in both context-stability and visual-consistency compared to SD and other inpainting algorithms.
Keywords:
Image Inpainting Context-Stability Visual-Consistency
Figure 1: Comparison on $1024^{2}$ image (depicted from MegaDepth [26]) between our proposed ASUKA and reconstruction-based (MAE [19]) as well as generative-based (MAE-FAR [7], MAT [25], Co-Mod [59], LaMa [45], Stable Diffusion [39]) inpainting models. MAE offers stable estimates, yet falls short in texture detail. GAN-based inpainting struggles with low fidelity. Powerful SD hallucinates random elements and suffers from mask-unmask visual inconsistency (color of clouds in the orange square). ASUKA achieves context-stable and visual-consistent inpainting.
@@ 1 Introduction
Figure 2: Object removal on $1024^{2}$ images. ASUKA enjoys context-stable (first to second rows) and visual-consistent (second to third rows) inpainting compared with SD.
@@
Image inpainting [5] fills masked areas of images while maintaining consistency with the unmasked regions. Traditional inpainting algorithms [5, 13, 18, 24, 40] often result in blurred synthesis when reconstructing masked regions, as indicated by [36]. The advent of Generative Adversarial Networks (GANs) has led to works [35, 27, 8, 45, 46, 59, 17, 21, 15] that utilize GANs to fill complex mask structures, achieving impressive inpainting results. However, they still struggle with general challenging inpainting cases, particularly in filling large holes.
@@
The recent Stable Diffusion Inpainting (SD) [39] model, with its large model capacity and extensive dataset training, acts as a versatile tool for image inpainting. However, we have observed a phenomenon of "context-instability" problem, where the model generates random, unreasonable elements to fill masked regions of images, as depicted in first to second rows in Fig. 2, as well as Fig. 1 and Fig. 8. This issue appears to come from the random masking strategy used to train SD. This strategy introduces training cases where foreground objects are completely masked and the SD is forced to fill masked regions with foreground objects. Consequently, SD will hallucinate unreasonable objects devoid of contextual information. Adjusting prompts may reduce this risk, but the best prompt is image-dependent, making it infeasible for practical applications.
Furthermore, the inpainted results of SD may suffer from "visual inconsistency" problem. This problem, less explored in academia but critical for real-world industrial applications, results in significant color discrepancies between inpainted and unmasked regions and exhibits smear-like traces in the image, as shown in the second to third rows in Fig. 2, as well as Fig. 1 and Fig. 8. Essentially, this visual-inconsistency comes from the misalignment between the pixel distributions of filled results and original images due to information loss in the Kullback-Leibler Variational Autoencoder (KL-VAE) used in SD, as illustrated in Fig. 5. To improve visual consistency, an inpainting-specialized VAE decoder is necessary to explicitly mitigate information loss based on available ground-truth information of unmasked images.
To efficiently tackle context-instability and visual inconsistency issues, we enhance the SD [39] with regression-based reconstruction and distribution-aligned generation. This results in an improved image inpainting model that avoids generating unreasonable elements in the masked region and reduces mask-unmask color inconsistency. The vanilla SD [39] adopts a KL-VAE to compress image into latent and perform diffusion in the latent space. We manipulate its generation and decoding processes for better context stability and visual consistency. i) We propose using the Masked Auto-Encoder (MAE)[19] as a prior to guide and stabilize the SD’s generation process. As shown in Fig. 2, MAE yields stable yet blurred results, while SD may produce implausible content despite its impressive generation capacity. By aligning MAE prior with SD objectives, we improve context-stability without compromising performance. ii) We redesign the KL-VAE decoder in SD to address color inconsistencies between masked and unmasked regions. The new decoder, acting as a local harmonization model conditioned on unmasked image pixels, significantly enhances visual consistency.
Formally, to enhance the context-stability and visual-consistency of image inpainting model, we present the Aligned Stable Inpainting with Unknown Areas prior (ASUKA) framework. i) To enhance context-stability, we replace the text-condition part of the SD with the MAE prior. We fine-tune the MAE to ensure adaptability to practical inpainting scenarios, as MAE is trained on uniformly distributed masks but inpainting usually contains large continuous masks. We train an alignment module to align MAE prior with SD using the standard SD objective while freezing the fine-tuned MAE and SD. ii) To address visual-consistency issues during decoding, we redesign the decoder of the KL-VAE in SD. We inject additional inputs of unmasked image pixels and 0-1 masks to the decoder. This new decoder is trained to perform local harmonization task, taking as inputs of inharmonious latent and taking as outputs the original harmonious image. We generate inharmonious latent via the proposed color and latent augmentation techniques to incorporate KL-VAE information loss and generation artifacts of SD. Then the decoder is trained to recover the original image from corrupted latent based on the ground-truth information of unmasked image. These steps collectively enable ASUKA to achieve more context-stable and visual-consistent inpainting results.
@@
To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, in addition to the benchmark dataset Places 2 [61], we further utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [10], Flickr-Landscape [28], MegaDept[26], and COCO 2014 [29]. This dataset thus covers four distinct domains—landscape, indoor, building, and background—making it diverse enough to serve as a benchmark for evaluation. Experiments on MISATO and Places 2, especially with large irregular masks, validate the efficacy of ASUKA.
@@
Contributions. ASUKA enhances image inpainting from context-stability and visual-consistency while leveraging the generation capacity of the frozen SD inpainting model. It achieves this through two main components: i) Context-Stable Alignment: ASUKA aligns the stable MAE prior with SD to provide a context-stable estimation of masked regions, replacing the text-condition part of SD with MAE prior. ii) Visual-Consistent Alignment: ASUKA re-formulates the decoding from SD latent to image as a local harmonization task, trains an inpainting-specialized decoder to align masked and unmasked regions during decoding and thus mitigates color inconsistencies caused by information loss from the SD.
@@ 2 Related Works
Image inpainting is a longstanding challenge to infill the missing regions with consistent pixels. Classical approaches, relying on patchmatch [14, 4, 56] and partial differential equation [5, 9, 6] are constrained by low-level image features and often fail to produce reasonable results for large holes. GAN-based [17] inpainting methods [36, 53, 59, 25, 7] have evolved with adaptive convolutions [30, 53, 54], attention mechanism [52, 51, 55, 23], and high-resolution generalization in frequency learning [45, 50, 12]. Methods like Co-Mod [59] prioritize adversarial learning and plausible generation to address the challenging ill-posed inpainting issue [25, 60], yielding robust and high-quality generations, even for large masks. However, they suffer from unstable results and uncontrollable hallucinations conditioned on different random latent codes. On the other hand, methods emphasizing higher reconstruction penalties [35, 45, 7] achieve more stable generations, but struggle with large masks, often resulting in blur artifacts. Recent diffusion models [41, 39] have rapidly developed and demonstrated impressive inpainting results. However, these models share similarities with GAN-based methods, learning to align distributions rather than achieving faithful pixel space reconstruction. As a consequence, diffusion-based approaches also suffer from context-instability issue.
Adapting Stable Diffusion. SD [39] is one of the most widely used diffusion models. SD incorporates a KL-VAE to learn a latent space preserving image semantics at a lower resolution, along with a U-Net for diffusion within this latent space. Various tuning and aligning methods have been developed to introduce additional conditions to SD, including image-inversion for text-guided image-to-image translation [34], textual-inversion for personalization [16], LoRA fine-tuning [22], and controlnet [57] to add other types of conditions. Our objective is to enhance context-stability while preserving the generation capacity of SD. Consequently, we refrain from fine-tuning the U-Net backbone of SD. The text-condition is not utilized in inpainting tasks. Textual-inversion [16] introduces learnable tokens in Clip text encoders to improve subject description. ObjectStitch [43] employs the CLIP visual feature to represent the subject. In our approach, we advocate removing the text-condition and instead guiding SD using the Masked Auto-Encoder (MAE) prior for masked region.
@@
Information loss in SD latent space. Although claimed only eliminates imperceptible details, the KL-VAE used by SD causes distortion in the reconstruction of images. See Fig. 5(a) for illustrative examples. OpenAI [2] proposes a larger decoder to improve the decoding quality of SD’s latent. Luo et al. [33] propose a frequency-augmented decoder to address the super-resolution case. Zhu et al. [62] propose to preserve unmasked regions during decoding. In this paper, we ensure the low-frequency color visual-consistency in the decoding process of SD.
@@
Masked Image-Modeling [3] (MIM) is an active research area in self-supervised learning. Typical MIM methods [3, 19, 49, 11] split images into visible and masked patches, learning to estimate masked patches from visible patches. Training targets for visible patches encompass pixel values [19], HOG features [47], and high-level semantic features [48]. While the primary objective of MIM is representation learning, its potential effectiveness in image generation is also noteworthy. Cao et al. [7] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies. In contrast, this paper uses MAE prior to enhance the context-stability of SD.
@@ 3 Methodology
Problem setup. Inpainting takes as inputs a masked image to complete with a mask to indicate the missing region. The target of inpainting is to fill the missing region based on the information of unmasked regions to generate high-fidelity images. In this paper, we focus on the standard inpainting task without utilizing other conditions. We adopt stable diffusion inpainting model [39] and focus on the general issues of inpainting models, i) context-instability issue: unstable and uncontrollable hallucinations, yielding random elements generated in the masked region; ii) visual-inconsistency issue: mask-unmask color inconsistency issue, yielding smear-like traces in the masked region.
@@
Figure 3: ASUKA tackles the context-instability issue by adopting the MAE to provide a stable prior for frozen SD to maintain the generation capacity while increasing context-stability. For the visual-inconsistency issue, ASUKA utilizes an inpainting-specialized decoder to achieve mask-unmask color consistency when decoding SD latent.
(a) Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
(b) Alignment module.
Figure 4: (a) Use MAE as initial value v.s. as condition. (b) Alignment module.
@@
Overview. The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Fig. 3. ASUKA adopts the pre-trained stable diffusion inpainting model (SD) [39]. Our target is to provide more context-stable and visual-consistent inpainting results while fully exploiting the generation capacity of SD. ASUKA includes i) a context-stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with SD and ii) a visual-consistent alignment to align ground-truth unmasked region with generated masked region during decoding. To this end, we freeze the U-Net of SD, while replacing the text-condition part with our proposed MAE prior to increase context-stability. To align the MAE prior to SD, we introduce an alignment module, which is trained via the training objective of SD. Additionally, to align masked and unmasked regions during decoding and resolve the problem of information loss from the KL-VAE decoder which causes mask-unmask color inconsistency, we train an inpainting-specialized decoder to decode the latent back to the image space for seamless visual-consistency. Combined together, ASUKA achieves more context-stable and visual-consistent inpainting.
3.1 Masked Auto-Encoder Prior
@@
MAE as context-stable prior. While recent generative models rely on random noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly. In contrast, MAE is known to provide a context-stable estimation of masked regions based purely on the unmasked regions. As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a context-stable prior for generative models to not generate new concepts. However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the SD to perform inpainting in image-to-image style, as illustrated in Fig. 3(a). To this end, we adopt the MAE to provide prior to stabilizing SD.
@@
Training MAE. The original MAE is trained to estimate random masks uniformly distributed in the image, while inpainting task usually contains large continuous masks. Inspired by Cao et al. [7], we fine-tune the MAE to inpainting masks. To adapt MAE for more practical inpainting scenarios, we construct a systematic masking strategy. The mask basis contains: object-shape mask, irregular mask, and regular mask. We collect object-shape masks from COCO [29] object segments. We use irregular masks from previous studies, including Co-Mod mask [59] and LaMa [45] mask. The regular masks contains rectangle and complement rectangle mask. To ensure generalization and coverage, for each mask we generate from mask basis with the probability of 50% object-shape, 40% irregular, and 10% regular. For object-shape mask basis, we combine it with irregular mask with the chance of 50%. This construction of mask style estimates the masks occurs in inpainting tasks, especially for the object removal and user-specified irregular masks. We control the mask ratio in the range of $[0.1,0.75]$ to follow the training scenario of MAE. For masks smaller than the ratio of 75%, we enlarge the mask ratio to 75% with randomly selected mask regions. This benefits ASUKA to tackle the large hole inpainting task.
3.2 Align MAE Prior with SD for Context-Stability

### v3

# Towards Enhanced Image Inpainting: Mitigating Unwanted Object Insertion and Preserving Color Consistency

@@ Abstract
Recent advances in image inpainting increasingly use generative models to handle large irregular masks. However, these models can create unrealistic inpainted images due to two main issues: (1) Unwanted object insertion: Even with unmasked areas as context, generative models may still generate arbitrary objects in the masked region that don’t align with the rest of the image. (2) Color inconsistency: Inpainted regions often have color shifts that causes a smeared appearance, reducing image quality. Retraining the generative model could help solve these issues, but it’s costly since state-of-the-art latent-based diffusion and rectified flow models require a three-stage training process: training a VAE, training a generative U-Net or transformer, and fine-tuning for inpainting. Instead, this paper proposes a post-processing approach, dubbed as ASUKA (Aligned Stable inpainting with UnKnown Areas prior), to improve inpainting models. To address unwanted object insertion, we leverage a Masked Auto-Encoder (MAE) for reconstruction-based priors. This mitigates object hallucination while maintaining the model’s generation capabilities. To address color inconsistency, we propose a specialized VAE decoder that treats latent-to-image decoding as a local harmonization task, significantly reducing color shifts for color-consistent inpainting. We validate ASUKA on SD 1.5 and FLUX inpainting variants with Places2 and MISATO, our proposed diverse collection of datasets. Results show that ASUKA mitigates object hallucination and improves color consistency over standard diffusion and rectified flow models and other inpainting methods.
footnotetext: First three authors contribute equally. Most parts of this work was done when Yikai was at Fudan. Yanwei Fu is the corresponding author.
@@ 1 Introduction
Figure 1: Image inpainting on $1024^{2}$ images. ASUKA solves two issues existed in current diffusion and rectified flow inpainting models: (1) Unwanted object insertion, where randomly elements that are not aligned with the unmasked region are generated; (2) Color-inconsistency: the color shift of the generated masked region, causing smear-like traces. ASUKA proposes a post-training procedure for these models, significantly mitigates object hallucination and improves color consistency of inpainted results.
@@
Image inpainting [6] fills masked areas of images while maintaining consistency with the unmasked regions. Traditional inpainting algorithms [6, 21, 30, 41, 69] often result in blurred synthesis when reconstructing masked regions [62]. The Generative Adversarial Networks (GANs) based models could fill complex mask structures, achieving impressive inpainting results [59, 46, 9, 75, 78, 99, 27, 33, 24]. However, they still struggle with general challenging inpainting cases, particularly in filling large holes. Recently, more powerful generative model like Stable Diffusion [67] and FLUX [40], with their large model capacity and extensive training dataset, act as versatile tools for image inpainting. These models mainly follow the latent generation pipeline, first encode the image into a small latent space, then train the inpainting model in this latent space.
However, these latent-based generative inpainting models still suffer from some issues, causing the inpainted image lacking fidelity. In particular:
@@
(1) The unwanted object insertion problem, where the model generates random, unreasonable elements to fill masked regions, as depicted in first to second rows in Fig. 1. This issue comes from the random masking strategy used to train generative models. This strategy introduces training cases where foreground objects are completely masked but the models are forced to fill masked regions with foreground objects. Consequently, these models will hallucinate unreasonable objects devoid of contextual information. Adjusting prompts may reduce this risk, but the best prompt is image-dependent, making it infeasible for practical applications.
(2) Furthermore, the inpainted results of latent inpainting models suffer from color inconsistency problem. This problem, less explored in academia but critical for real-world applications, results in color discrepancies between inpainted and unmasked regions, including mismatches of brightness, saturation, luminance and hue, and exhibits smear-like traces in the image, as shown in the second to third rows in Fig. 1. Essentially, this color-inconsistency comes from the misalignment between the pixel distributions of filled results and original images due to imperfect latent generative model and VAE, as illustrated in Fig. 4. Notably, this issue is not a big problem for generation, given that the whole image is generated, and the color shift is consistent across pixels. However, this issue is important for inpainting tasks, as we have ground-truth pixels for unmasked regions. When we replace the unmasked regions of the generated image with the ground-truth pixels, the color inconsistency largely influence the fidelity of the image. This issue may be solved by training a better VAE and explicitly enforce the color consistency. However, training or fine-tuning the VAE encoder introduces the subsequent fine-tuning of the latent generative models to match the new latent space, which is costly. In this paper, we propose to freeze the VAE encoder and the latent generative models, while fine-tuning the VAE decoder to improve color consistency. Specifically, we reformulate the decoding from latent to image as a local harmonization task, explicitly reduce the color inconsistency.
Formally, to mitigate object hallucination and enhance the color-consistency of image inpainting models, we present the Aligned Stable inpainting with UnKnown Areas prior (ASUKA) framework. ASUKA enhances the latent inpainting models with regression-based reconstruction and distribution-aligned generation. This results in improved image inpainting models that avoid generating unreasonable elements in the masked region and reduces mask-unmask color inconsistency. The stable diffusion models [67] and the rectified flow models [40] adopt a VAE to compress image into latent and perform inpainting in the latent space. We manipulate their generation and decoding processes to reduce object hallucination and improve color consistency.
We propose using the Masked Auto-Encoder (MAE)[31] as a prior to guide and stabilize the generation process. As shown in Fig. 1, MAE yields stable yet blurred results, while generative models may produce implausible content despite their impressive generation capacity. By aligning MAE prior with latent generative models, we reduce object hallucination without damaging performance.
We redesign the VAE decoder to address color inconsistencies between masked and unmasked regions by acting as a local harmonization model conditioned on unmasked image pixels. Our decoder can be used as a plug-and-play module to improve general inpainting models, such as text-guided inpainting.
@@
These steps collectively enable ASUKA to achieve less object hallucination and more color-consistent inpainting results. We adopt ASUKA on two typical inpaitning models, Stable Diffusion v1.5 [67] and FLUX [40], to validate the generalization ability of ASUKA on different generation architectures. To evaluate the effectiveness of inpainting algorithms across various scenarios and mask shapes, in addition to the benchmark dataset Places 2 [101], we further utilize an evaluation dataset named MISATO, which selects representative testing images from Matterport3D [13], Flickr-Landscape [47], MegaDept[45], and COCO 2014 [48]. This dataset covers four distinct domains—landscape, indoor, building, and background—making it diverse to serve as a benchmark for evaluation. Experiments on MISATO and Places 2 with large irregular masks validate the efficacy of ASUKA.
Contributions
@@
ASUKA enhances image inpainting with color-consistency and mitigate object hallucination while leveraging the generation capacity of the frozen inpainting model. It achieves this through two main components: (1) Context-Stable Alignment: ASUKA aligns the stable MAE prior with generative models to provide a context-stable estimation of masked regions, replacing the text-condition with MAE prior. (2) Color-Consistent Alignment: ASUKA re-formulates the decoding from latent to image as a local harmonization task, trains an inpainting-specialized decoder to align masked and unmasked regions during decoding and thus mitigates color inconsistencies.
@@ 2 Related Works
Image inpainting
is the task of filling missing image regions with consistent pixels. Traditional methods using patch matching [22, 5, 95] or differential equations [6, 12, 7] focus on low-level features and often struggle with large gaps. GAN [27]-based inpainting [62, 91, 99, 44, 10] introduces adaptive convolutions [50, 91, 93], attention [90, 89, 92, 39], and frequency-based learning for high-resolution results [75, 87, 17]. Methods like Co-Mod [99] address the challenging ill-posed inpainting issue [44, 100] and improve realism but may produce unstable outputs or unwanted artifacts due to random latent variables. Techniques with higher reconstruction penalties [59, 75, 10] offer more stability but can appear blurry on larger missing areas. Recent diffusion models [70, 67, 3] and rectified flow models [25, 40] achieve impressive results yet share GANs’ limitation of learning distributions over exact pixel alignment, which leads to unwanted object insertion.
Adapting latent generative models
Latent diffusion models (LDMs) [67] and rectified flow models [25, 40] are popular due to their ability to encode image semantics at lower resolutions by combining a VAE to learn a latent space and a generative model within this space. Various methods have been developed to introduce new conditions to these models, such as image-inversion for text-guided image translation [56], textual-inversion for personalization [26], LoRA fine-tuning [34], and controlnet [96] to add diverse conditions. Our goal is to mitgate object hallucination while preserving generation quality, so we avoid fine-tuning the generative backbone. For inpainting, we remove the text condition and instead guide the generation using a Masked Auto-Encoder [31] prior for masked regions.
Information loss in latent inpainting models
@@
Although claimed only eliminates imperceptible details, the VAE used by diffusion and rectified flow models causes distortion in the reconstruction of images. In addition, the gap between generated latent and real latent also causes the color inconsistency. See Fig. 4 for illustrative examples. OpenAI [61] proposes a larger decoder to improve the decoding quality of SD’s latent. Luo et. al. [55] propose a frequency-augmented decoder to address the super-resolution case. Zhu et. al. [103] propose to preserve unmasked regions during decoding. In this paper, we ensure the low-frequency color color-consistency in the decoding process.
Masked Image-Modeling [4]
@@
(MIM) is an active research area in self-supervised learning. Typical MIM methods [4, 31, 86, 14] split images into visible and masked patches, learning to estimate masked patches from visible patches. Training targets for visible patches encompass pixel values [31], HOG features [82], and high-level semantic features [83]. While the primary objective of MIM is representation learning, its potential effectiveness in image generation is also noteworthy. Cao et. al. [10] adopts MAE features and attention scores to assist the convolutional inpainting model better in handling long-distance dependencies. In contrast, this paper uses MAE prior to enhance the context-stability of diffusion and rectified flow models.
Image harmonization
aims to blend a foreground object with a background image while keeping the final result realistic and visually consistent [76]. This task is often treated as an image translation problem [102, 19, 28, 20, 29, 60, 79, 51, 57, 66]. Similarly, our work addresses color inconsistency issues in latent generative models. However, unlike standard image harmonization, where inconsistencies arise from combining images from different sources and thus different real image distributions, color inconsistencies in latent generative models stem from imperfections in the VAE and the generative model itself.
Object insertion and removal
are two opposite tasks in image inpainting. Object insertion focuses on adding foreground objects to the image using various methods, such as shape-guided masks [94, 85], text prompts [80, 85, 8, 16], learnable prompts [81, 104, 16], extra network modules [15, 35], or reference images of objects [71], etc. Some studies also explore completing partial objects using reference images [11] or learnable prompts [81]. Object removal, on the other hand, aims to erase unwanted objects from an image. Common approaches include attention reweighting [42] and learnable prompts [81, 104]. These techniques can help create new datasets [23]. On the other hand, creating new datasets can also benefit these tasks [84]. While most research focuses on designing better inpainting models, our work takes a different approach. We analyze a fundamental problem with latent generative models: they often introduce unwanted objects in the inpainting area. We also propose solutions to address this issue.
@@ 3 Methodology
Inpainting takes as inputs a masked image to complete with a mask to indicate the missing region. The target of inpainting is to fill the missing region based on the information of unmasked regions to generate high-fidelity images. In this paper, we focus on the standard inpainting task without utilizing other conditions. We focus on the general issues of inpainting models, (1) unwanted object insertion: unstable and uncontrollable hallucinations, yielding random elements generated in the masked region; (2) color-inconsistency: mask-unmask color inconsistency issue, yielding smear-like traces in the masked region.
@@
Figure 2: ASUKA tackles the unwanted object insertion issue by adopting the MAE to provide a stable prior for frozen latent generative models to maintain the generation capacity while mitigating object hallucination. For the color-inconsistency issue, ASUKA utilizes an inpainting-specialized decoder to achieve mask-unmask color consistency when decoding latent.
We evaluate our proposed solution on two inpainting models: the Stable Diffusion v1.5 inpainting model (SD) [67] and the Control-Net fine-tuned FLUX inpainting model (FLUX) [2]. We provide a brief introduction of these models in the appendix. We will demonstrate that our ASUKA effectively improves unwanted object mitigation and color consistency of these models.
Overview
@@
The framework of the proposed Aligned Stable inpainting with Unknown Areas prior (ASUKA) is illustrated in Fig. 2(a). ASUKA adopts the pre-trained latent inpainting models. Our target is to mitigate object hallucination and provide more color-consistent inpainting results while fully exploiting the generation capacity of frozen models. ASUKA includes (1) a context-stable alignment to align stable Masked Auto-Encoder (MAE) prior for masked region with generative models and (2) a color-consistent alignment to align ground-truth unmasked region with generated masked region during decoding. To this end, we freeze the latent generative models, while replacing the text-condition part with our proposed MAE prior to mitigate object hallucination. To align the MAE prior to generative models, we introduce an alignment module, trained via the training objective of generative models. Additionally, to align masked and unmasked regions during decoding and resolve the information loss issue from VAE decoder and generative model which causes mask-unmask color inconsistency, we train an inpainting-specialized decoder to decode the latent back to the image space for seamless color-consistency. Combined together, ASUKA achieves less object hallucination and more color-consistent inpainting.
3.1 Mitigate Object Hallucination via Stable Prior
3.1.1 Masked Auto-Encoder Prior
Context-stable prior
@@ 3.1.1 Masked Auto-Encoder Prior
While recent generative models rely on random noise to provide more diverse generation results, it leads to the generation of random objects unexpectedly. Some inpainting models also utilize the reconstruction loss to reconstruct the masked region, but they also incorporate other types of losses like perceptual-loss [75] which implicitly reduces the stability. In contrast, MAE is known to provide a context-stable estimation of masked regions based purely on the unmasked regions. In this paper, we utilize MAE to produce the stable prior such that the improvement of inpainted result can be explicitly attributed to the improvement of mitigating object hallucination.
MAE as context-stable prior
As MAE is trained on the L2 reconstruction loss, we can regard the estimation of MAE as a mean estimation, which can be utilized to provide a context-stable prior for generative models to not generate new concepts. However, MAE itself results in average and blur generations and cannot reconstruct detailed textures of the masked region, and works poorly if we use MAE prior as the initial values for the inpainting models to inpaint in image-to-image style, as in Fig. 3. To this end, we adopt the MAE to provide prior to stabilizing diffusion models.
Train MAE
@@
The original MAE is trained to estimate random masks uniformly distributed in the image, while inpainting task usually contains large continuous masks. Inspired by Cao et. al. [10], we fine-tune the MAE to inpainting masks. To adapt MAE for more practical inpainting scenarios, we construct a systematic masking strategy. The mask basis contains: object-shape mask, irregular mask, and regular mask. We collect object-shape masks from COCO [48] object segments. We use irregular masks from previous studies, including Co-Mod mask [99] and LaMa [75] mask. The regular masks contains rectangle and complement rectangle mask. To ensure generalization and coverage, for each mask we generate from mask basis with the probability of 50% object-shape, 40% irregular, and 10% regular. For object-shape mask basis, we combine it with irregular mask with the chance of 50%. This construction of mask style estimates the masks occurs in inpainting tasks, especially for the object removal and user-specified irregular masks. We control the mask ratio in the range of $[0.1,0.75]$ to follow the training scenario of MAE. For masks smaller than the ratio of 75%, we enlarge the mask ratio to 75% with randomly selected mask regions. This benefits ASUKA to tackle the large hole inpainting task.
Figure 3: Use MAE prior for image-to-image translation (start from 80% noise rate) via SD achieves poor inpainting results.
3.1.2 Align MAE Prior with Generator
Replace text-condition with MAE prior